# S2AFormer: Strip Self-Attention for Efficient Vision Transformer

### v1

@@ III-C Complexity Analysis
To demonstrate the efficiency of our proposed SSA, we compare its computational overhead with standard self-attention [2]. Given an input feature $x\in\mathbb{R}^{N\times d}$, we analyze the computational cost, measured in FLOPs, as follows.
The projection step introduces the complexity of
$\mathcal{O}(\text{proj})=Nd(d_{q}+d_{k}+d_{v}),$ (8)
where $d_{q}$, $d_{k}$, and $d_{v}$ represent the dimensions of the query, key, and value, respectively. Calculating the attention weights involves the dot product of $Q$ and $K^{T}$, which incurs a complexity of
$\mathcal{O}(Q\cdot K^{T})=N^{2}\cdot d_{q}.$ (9)
Subsequently, applying the attention weights to $V$ requires
$\mathcal{O}(\text{Attn}\cdot V)=N^{2}\cdot d_{v}.$ (10)
Combining these, the total computational cost of Multi-Head Self-Attention (MHSA) is given by
@@
In the case of our SSA, a spatial reduction is introduced, with $N_{s}=\frac{N}{k^{2}}$, where $k$ represents the kernel size of the spatial reduction convolution. The projection phase, incorporating the effects of spatial reduction, has a complexity of
$\Omega(\text{proj})=Nd\left(h+\frac{h}{k^{2}}+\frac{d}{k^{2}}\right),$ (13)
where $h\ll d$ is the number of attention heads. The computation of attention weights involves
$\Omega(Q\cdot K^{T})=N\cdot N_{s}=\frac{N^{2}}{k^{2}}.$ (14)
Similarly, multiplying the attention weights by $V$ results in a complexity of
$\Omega(\text{Attn}\cdot V)=NN_{s}\cdot d=\frac{N^{2}d}{k^{2}}.$ (15)
@@
$\mathcal{O}(\text{SSA})=\frac{1}{k^{2}}Nd^{2}+\frac{1+d}{k^{2}}N^{2}+\left(1+\frac{1}{k^{2}}\right)Ndh.$ (16)
@@
$\mathcal{O}(\text{SSA})\ll\frac{1}{{{k^{2}}}}N{d^{2}}+\frac{{1+d}}{{{k^{2}}}}{N^{2}}+\left({1+\frac{1}{{{k^{2}}}}}\right)N{d^{2}}.$ (17)
Let $\mathcal{O}(\text{A})=\frac{1}{{{k^{2}}}}N{d^{2}}+\frac{{1+d}}{{{k^{2}}}}{N^{2}}+\left({1+\frac{1}{{{k^{2}}}}}\right)N{d^{2}}$, the complexity can be further simplified as
$\mathcal{O}(\text{A})$ $=\left(1+\frac{2}{k^{2}}\right)Nd^{2}+\frac{1+d}{k^{2}}N^{2}$ (18)
$<3Nd^{2}+2dN^{2},$
@@
TABLE III: Comparison of image classification performance on the ImageNet-1K dataset [18].
@@
MobileFormer-52M [56] Conv+Attn $224^{2}$ 3.5 0.052 68.7
MobileViT-XXS [55] Conv+Attn $256^{2}$ 1.3 0.364 69.0
MobileNet-v1 [65] Conv $224^{2}$ 4.2 0.575 70.6
EfficientViT-M2 [11] Conv+Attn $224^{2}$ 4.2 0.201 70.8
EdgeNeXt-XXS [44] Conv+Attn $256^{2}$ 1.3 0.261 71.2
EMO-1M [13] Conv $224^{2}$ 1.3 0.26 71.5
FasterNet-T0 [66] Conv+Attn $224^{2}$ 3.9 0.34 71.9
MobileNet-v3 [65] Conv $256^{2}$ 1.4 0.481 72.3
SHViT-S1 [15] Conv+Attn $224^{2}$ 6.3 0.24 72.8
Vim-Ti [67] Conv+SSM $224^{2}$ 7 1.5 73.1
S2AFormer-mini (Ours) Conv+Attn $224^{2}$ 5.02 0.43 75.1
S2AFormer-mini (Ours) Conv+Attn $256^{2}$ 5.02 0.56 75.7
MobileMamba-T2 [68] Conv+SSM $192^{2}$ 8.8 0.26 73.6
EdgeViT-XXS [12] Conv+Attn $224^{2}$ 4.1 0.6 74.4
DeiT-T [5] Attn $224^{2}$ 5.9 1.2 74.5
MobileViT-XS [55] Conv+Attn $224^{2}$ 2.3 0.706 74.8
SHViT-S2 [15] Conv+Attn $224^{2}$ 11.4 0.37 75.2
InceptionNeXt-A [17] Conv $224^{2}$ 4.2 0.51 75.3
FastViT-T8 [14] Conv+Attn $256^{2}$ 3.6 0.7 75.6
LocalVim-T [69] Conv+SSM $224^{2}$ 8 1.5 76.2
EfficientVMamba-T [70] Conv+SSM $224^{2}$ 6.0 0.8 76.5
SHViT-S3 [15] Conv+Attn $224^{2}$ 14.2 0.60 77.4
S2AFormer-T (Ours) Conv+Attn $224^{2}$ 5.80 0.655 77.7
S2AFormer-T (Ours) Conv+Attn $256^{2}$ 5.80 0.855 78.3
EfficientViT-M5 [11] Conv+Attn $224^{2}$ 12.4 0.522 77.1
MobileViG-S [71] Conv+GNN $224^{2}$ 7.2 1.0 78.2
EdgeViT-XS [12] Conv+Attn $224^{2}$ 6.7 1.1 77.5
MobileFormer-294M [56] Conv+Attn $224^{2}$ 11.4 0.294 77.9
SwiftFormer-S [42] Conv+Attn $224^{2}$ 6.1 1.0 78.5
PoolFormer-S12 [30] Pool $224^{2}$ 12 1.9 77.2
PlainMamba-L1 [72] Conv+SSM $224^{2}$ 7 3.0 77.9
EdgeNeXt-S [44] Conv+Attn $224^{2}$ 5.6 0.965 78.8
RepViT-M0.9 [31] Conv $224^{2}$ 5.1 0.8 78.7
EfficientMod-xs [48] Conv $224^{2}$ 6.6 0.8 78.3
EfficientVMamba-S [70] Conv+SSM $224^{2}$ 11 1.3 78.7
MobileViT-S [55] Conv+Attn $224^{2}$ 5.6 2.01 78.4
StarNet-S4 [73] Conv $224^{2}$ 7.5 1.075 78.4
S2AFormer-XS (Ours) Conv+Attn $224^{2}$ 6.54 0.786 78.9
S2AFormer-XS (Ours) Conv+Attn $256^{2}$ 6.54 1.030 79.3
FastViT-T12 [14] Conv+Attn $256^{2}$ 6.8 1.4 79.1
EfficientFormer-L1 [43] Conv+Attn $224^{2}$ 12.3 1.3 79.2
EdgeNeXt-S [44] Conv+Attn $256^{2}$ 5.6 1.30 79.4
SHViT-S4 [15] Conv+Attn $256^{2}$ 16.5 0.99 79.4
PoolFormer-S24 [30] Pool $224^{2}$ 21 3.5 80.3
Vim-S [67] Conv+SSM $224^{2}$ 26 5.1 80.3
MobileViG-M [71] Conv+GNN $224^{2}$ 14.0 1.5 80.6
SwiftFormer-L1 [42] Conv+Attn $224^{2}$ 12.1 1.62 80.9
EfficientViT-M5 [11] Conv+Attn $512^{2}$ 12.4 2.67 80.8
FastViT-SA12 [14] Conv+Attn $256^{2}$ 10.9 1.9 80.6
RepViT-M1.1 [31] Conv $224^{2}$ 8.2 1.3 80.7
EdgeViT-S [12] Conv+Attn $224^{2}$ 11.1 1.9 81.0
EfficientMod-s [48] Conv $224^{2}$ 12.9 1.4 81.0
S2AFormer-S (Ours) Conv+Attn $224^{2}$ 10.69 1.38 80.8
S2AFormer-S (Ours) Conv+Attn $256^{2}$ 10.69 1.80 81.3
RMT-T [74] Conv $224^{2}$ 14.0 2.5 82.4
TransNeXt-Micro [75] Conv+Attn $224^{2}$ 12.8 2.7 82.5
EfficientFormer-L3 [43] Conv+Attn $224^{2}$ 31.3 3.9 82.4
MobileViG-Ti [71] Conv+GNN $224^{2}$ 26.7 2.8 82.6
RepViT-M1.5 [31] Conv $224^{2}$ 14.0 2.3 82.3
RepNeXt-M4 [60] Conv $224^{2}$ 13.3 2.3 81.2
CAS-ViT-M [61] Conv $224^{2}$ 12.42 1.887 81.4
SHViT-S4 [15] Conv+Attn $512^{2}$ 16.5 3.97 82.0
VMamba-T [76] Conv+SSM $224^{2}$ 22 5.6 82.2
MambaOut-Tiny [77] Conv $224^{2}$ 27 4.5 82.7
EfficientVMamba-B [70] Conv+SSM $224^{2}$ 33 4.0 81.8
LocalVim-S [69] Conv+SSM $224^{2}$ 28 4.8 81.2
InceptionNeXt-T [17] Conv $224^{2}$ 28 4.2 82.3
PeLK-T [78] Conv $224^{2}$ 29 5.6 82.6
S2AFormer-M (Ours) Conv+Attn $224^{2}$ 24.87 4.12 82.3
S2AFormer-M (Ours) Conv+Attn $256^{2}$ 24.87 5.38 82.7
TABLE IV: Comparison of semantic segmentation performance on the ADE20K dataset [79].
Model Semantic FPN

### v2

@@ III-A Overall Architecture
III-B Hybrid Perception Block (HPB)
The Hybrid Perception Block (HPB) framework, shown in Fig. 3 (c), consists of two main components: the Strip Self-Attention (SSA) module in Fig. 3 (d) and the Local Interaction Module (LIM) in Fig. 3 (e). The overall operation is defined as follows:
$f_{Conv}=DWConv\left(x\right)+x,$ (1)
$f_{ssa}=SSA\left({LN\left(f_{Conv}\right)}\right)+f_{Conv},$ (2)
$f_{lim}=LIM\left({LN\left(f_{ssa}\right)}\right)+f_{ssa}.$ (3)
Finally, the MLP is conducted, as:
$f_{mlp}=MLP\left({LN\left(f_{lim}\right)}\right)+f_{lim}.$ (4)
TABLE III: Complexity comparison between Multi-Head Self-Attention (MHSA) and Strip Self-Attention (SSA).
Stage Operation Computation Complexity
MHSA SSA
Linear Projection $Q=xW_{Q}\in\mathbb{R}^{N\times d_{q}}$ $\mathcal{O}(Nd(d_{q}+d_{k}+d_{v}))$ $\mathcal{O}(Nd(h+\frac{h}{k^{2}}+\frac{d_{v}}{k^{2}}))$
$K=xW_{K}\in\mathbb{R}^{N\times d_{k}}$
$V=xW_{V}\in\mathbb{R}^{N\times d_{v}}$
Attention Scores $QK^{T}\in\mathbb{R}^{N\times N}$ $\mathcal{O}(N^{2}d_{q})$ $\mathcal{O}(\frac{N^{2}}{k^{2}}\cdot h)$
Weighted Sum $Attn\cdot V\in\mathbb{R}^{N\times d_{v}}$ $\mathcal{O}(N^{2}d_{v})$ $\mathcal{O}(\frac{N^{2}}{k^{2}}d_{v})$
Total Complexity $h\ll d_{q}=d_{k}=d_{v}=d$ $\mathcal{O}(3Nd^{2}+2dN^{2})$ $\mathcal{O}(\frac{1}{k^{2}}Nd^{2}+\frac{d+h}{k^{2}}N^{2}+(1+\frac{1}{k^{2}})Ndh)$
$\ll(1+\frac{2}{k^{2}})Nd^{2}+\frac{2}{k^{2}}dN^{2}$
$3Nd^{2}+2dN^{2}$
(a) ResNet [krizhevsky2012imagenet]
(b) CAS-ViT [zhang2024cas]
(c) EMO [zhang2023rethinking]
(d) Vanilla ViT [dosovitskiy2020image]
(e) SwiftFormer [shaker2023swiftformer]
(f) Ours
Figure 5: Visualization of effective respective fields (ERFs) across different models. Convolution-based models (a), (b), and (c) exhibit highly localized receptive fields, while Vanilla ViT (d) distributes attention broadly across all spatial positions. Model (e) demonstrates more limited receptive fields compared to ours. In contrast, our method (f) achieves a balanced pattern—effectively capturing key local regions and progressively expanding outward in a strip-like manner.
TABLE IV: Comparison of image classification performance on the ImageNet-1K dataset [deng2009imagenet]. Mixer refers to the token mixer operation, Res. indicates the input resolution of training, and #Para. corresponds to the total number of parameters.
[…]
MobileFormer-52M [chen2022mobile] Conv+Attn $224^{2}$ 3.50 0.052 68.7
MobileViT-XXS [mehta2021mobilevit] Conv+Attn $256^{2}$ 1.30 0.364 69.0
MobileViT-v2-0.5 [mehta2022separable] Attn $256^{2}$ 1.40 0.500 70.2
EfficientViT-M2 [liu2023efficientvit] Conv+Attn $224^{2}$ 4.20 0.201 70.8
EdgeNeXt-XXS [maaz2022edgenext] Conv+Attn $256^{2}$ 1.30 0.261 71.2
EMO-1M [zhang2023rethinking] Conv $224^{2}$ 1.30 0.260 71.5
FasterNet-T0 [chen2023run] Conv+Attn $224^{2}$ 3.90 0.340 71.9
MobileNet-v3 [howard2017mobilenets] Conv $256^{2}$ 1.40 0.481 72.3
EMOv2-1M [zhang2025emov2] Conv $224^{2}$ 1.40 0.285 72.3
SHViT-S1 [yun2024shvit] Conv+Attn $224^{2}$ 6.30 0.240 72.8
Vim-Ti [zhu2024vision] Conv+SSM $224^{2}$ 7.00 1.500 73.1
MobileMamba-T2 [he2025mobilemamba] Conv+SSM $192^{2}$ 8.80 0.255 73.6
iFormer-T [zheng2025iformer] Conv+Attn $224^{2}$ 2.90 0.530 74.1
LSNet-T [wang2025lsnet] Conv+Attn $224^{2}$ 11.40 0.300 74.9
VRWKV-T [duan2024vrwkv] Conv+Attn $224^{2}$ 6.20 1.200 75.1
S2AFormer-mini (Ours) Conv+Attn $224^{2}$ 5.02 0.430 75.1
S2AFormer-mini (Ours) Conv+Attn $256^{2}$ 5.02 0.560 75.7
EdgeViT-XXS [chen2022edgevit] Conv+Attn $224^{2}$ 4.10 0.600 74.4
DeiT-T [touvron2021training] Attn $224^{2}$ 5.90 1.200 74.5
MobileViT-XS [mehta2021mobilevit] Conv+Attn $224^{2}$ 2.30 0.706 74.8
SHViT-S2 [yun2024shvit] Conv+Attn $224^{2}$ 11.40 0.370 75.2
InceptionNeXt-A [yu2024inceptionnext] Conv $224^{2}$ 4.20 0.510 75.3
FastViT-T8 [vasu2023fastvit] Conv+Attn $256^{2}$ 3.60 0.700 75.6
EMOv2-2M Conv $224^{2}$ 2.30 0.487 75.8
MobileMamba-T4 [he2025mobilemamba] Conv+SSM $192^{2}$ 14.20 0.413 76.1
EfficientVMamba-T [pei2024efficientvmamba] Conv+SSM $224^{2}$ 6.00 0.800 76.5
EfficientViT-M5 [liu2023efficientvit] Conv+Attn $224^{2}$ 12.40 0.522 77.1
PoolFormer-S12 [yu2022metaformer] Pool $224^{2}$ 12.00 1.900 77.2
SHViT-S3 [yun2024shvit] Conv+Attn $224^{2}$ 14.20 0.600 77.4
LSNet-S [wang2025lsnet] Conv+Attn $224^{2}$ 16.10 0.500 77.8
FAT-B0 [fan2023lightweight] Conv+Attn $224^{2}$ 4.50 0.700 77.6
S2AFormer-T (Ours) Conv+Attn $224^{2}$ 5.80 0.655 77.7
S2AFormer-T (Ours) Conv+Attn $256^{2}$ 5.80 0.855 78.3
MobileViG-S [munir2023mobilevig] Conv+GNN $224^{2}$ 7.20 1.000 78.2
EdgeViT-XS [chen2022edgevit] Conv+Attn $224^{2}$ 6.70 1.100 77.5
MobileFormer-294M [chen2022mobile] Conv+Attn $224^{2}$ 11.40 0.294 77.9
SwiftFormer-S [shaker2023swiftformer] Conv+Attn $224^{2}$ 6.10 1.000 78.5
MobileViT-v2-1.0 [mehta2022separable] Attn $256^{2}$ 4.90 1.800 78.1
EdgeNeXt-S [maaz2022edgenext] Conv+Attn $224^{2}$ 5.60 0.965 78.8
MobileMamba-S6 [he2025mobilemamba] Conv+SSM $224^{2}$ 15.00 0.652 78.0
RepViT-M0.9 [wang2024repvit] Conv $224^{2}$ 5.10 0.800 78.7
FastViT-T12 [vasu2023fastvit] Conv+Attn $256^{2}$ 6.80 1.400 79.1
EfficientFormer-L1 [li2022efficientformer] Conv+Attn $224^{2}$ 12.30 1.300 79.2
EfficientMod-xs [ma2024efficient] Conv $224^{2}$ 6.60 0.800 78.3
EfficientVMamba-S [pei2024efficientvmamba] Conv+SSM $224^{2}$ 11.00 1.300 78.7
MobileViT-S [mehta2021mobilevit] Conv+Attn $224^{2}$ 5.60 2.010 78.4
StarNet-S4 [ma2024rewrite] Conv $224^{2}$ 7.50 1.075 78.4
S2AFormer-XS (Ours) Conv+Attn $224^{2}$ 6.54 0.786 78.9
S2AFormer-XS (Ours) Conv+Attn $256^{2}$ 6.54 1.030 79.3
[…]
EdgeNeXt-S [maaz2022edgenext] Conv+Attn $256^{2}$ 5.60 1.300 79.4
SHViT-S4 [yun2024shvit] Conv+Attn $256^{2}$ 16.50 0.990 79.4
MobileMamba-B1 [he2025mobilemamba] Conv+SSM $256^{2}$ 17.10 1.080 79.9
PoolFormer-S24 [yu2022metaformer] Pool $224^{2}$ 21.00 3.500 80.3
FAT-B1 [fan2023lightweight] Conv+Attn $224^{2}$ 7.80 1.200 80.1
VRWKV-S [duan2024vrwkv] Conv+Attn $224^{2}$ 23.80 4.600 80.1
Vim-S [zhu2024vision] Conv+SSM $224^{2}$ 26.00 5.100 80.3
LSNet-B [wang2025lsnet] Conv+Attn $224^{2}$ 23.20 1.300 80.3
MobileViG-M [munir2023mobilevig] Conv+GNN $224^{2}$ 14.00 1.500 80.6
SwiftFormer-L1 [shaker2023swiftformer] Conv+Attn $224^{2}$ 12.10 1.620 80.9
EfficientViT-M5 [liu2023efficientvit] Conv+Attn $512^{2}$ 12.40 2.670 80.8
FastViT-SA12 [vasu2023fastvit] Conv+Attn $256^{2}$ 10.90 1.900 80.6
RepViT-M1.1 [wang2024repvit] Conv $224^{2}$ 8.20 1.300 80.7
EdgeViT-S [chen2022edgevit] Conv+Attn $224^{2}$ 11.10 1.900 81.0
EfficientMod-s [ma2024efficient] Conv $224^{2}$ 12.90 1.400 81.0
S2AFormer-S (Ours) Conv+Attn $224^{2}$ 10.69 1.380 80.8
S2AFormer-S (Ours) Conv+Attn $256^{2}$ 10.69 1.800 81.3
RMT-T [fan2024rmt] Conv $224^{2}$ 14.00 2.500 82.4
TransNeXt-Micro [shi2024transnext] Conv+Attn $224^{2}$ 12.80 2.700 82.5
EfficientFormer-L3 [li2022efficientformer] Conv+Attn $224^{2}$ 31.30 3.900 82.4
MobileViG-Ti [munir2023mobilevig] Conv+GNN $224^{2}$ 26.70 2.800 82.6
RepViT-M1.5 [wang2024repvit] Conv $224^{2}$ 14.00 2.300 82.3
CAS-ViT-M [zhang2024cas] Conv $224^{2}$ 12.42 1.887 81.4
MobileMamba-B2 [he2025mobilemamba] Conv+SSM $384^{2}$ 17.10 2.427 81.6
VRWKV-B [duan2024vrwkv] Conv+Attn $224^{2}$ 93.70 18.200 82.0
SHViT-S4 [yun2024shvit] Conv+Attn $512^{2}$ 16.50 3.970 82.0
MobileMamba-B4 [he2025mobilemamba] Conv+SSM $512^{2}$ 17.10 4.313 82.5
MambaOut-Tiny [yu2024mambaout] Conv $224^{2}$ 27.00 4.500 82.7
EfficientVMamba-B [pei2024efficientvmamba] Conv+SSM $224^{2}$ 33.00 4.000 81.8
InceptionNeXt-T [yu2024inceptionnext] Conv $224^{2}$ 28.00 4.200 82.3
PeLK-T [chen2024pelk] Conv $224^{2}$ 29.00 5.600 82.6
S2AFormer-M (Ours) Conv+Attn $224^{2}$ 24.87 4.120 82.3
S2AFormer-M (Ours) Conv+Attn $256^{2}$ 24.87 5.380 82.7
ConvNeXt-B [liu2022convnet] Conv $224^{2}$ 89.00 15.400 83.8
Swin-B [liu2021swin] Attn $224^{2}$ 88.00 15.400 83.5
EfficientFormer-L7 [li2022efficientformer] Conv+Attn $224^{2}$ 82.10 10.200 83.3
PlainMamba-L3 [yang2024plainmamba] Conv+SSM $224^{2}$ 50.00 14.400 82.3
Focal-Base [yang2021focal] Attn $224^{2}$ 89.80 16.000 83.8
GroupMixFormer-B [ge2023advancing] Attn $224^{2}$ 45.80 17.600 84.7
CAFormer-M36 [yu2024metaformer] Attn $224^{2}$ 56.00 13.200 85.2
VRWKV-L [duan2024vrwkv] Conv+Attn $84^{2} 33$4.90 189.500 86.0
RAVLT-L [fan2025breaking] Attn $224^{2}$ 95.00 16.000 85.8
XCiT-M24/16 [ali2021xcit] Attn $224^{2}$ 84.00 16.200 84.3
InceptionNeXt-B [yu2024inceptionnext] Conv $224^{2}$ 87.00 14.900 84.0
PoolFormer-M48 [yu2022metaformer] Conv $224^{2}$ 73.00 11.600 82.5
Conv2Former-B [hou2024conv2former] Conv+Attn $224^{2}$ 90.00 15.900 84.4
TransNeXt-Base [shi2024transnext] Conv+Attn $224^{2}$ 89.70 18.400 84.8
S2AFormer-L (Ours) Conv+Attn $224^{2}$ 76.58 12.530 85.6
S2AFormer-L (Ours) Conv+Attn $256^{2}$ 76.58 16.370 86.0
III-B1 Strip Self-Attention (SSA)
In the standard self-attention mechanism [dosovitskiy2020image], the input $x\in{\mathbb{R}^{N\times C}}$, where $N=HW$, is linearly projected into the query $Q\in{\mathbb{R}^{N\times d_{q}}}$, the key $K\in{\mathbb{R}^{N\times d_{k}}}$, and the value $V\in{\mathbb{R}^{N\times d_{v}}}$. Subsequently, the self-attention mechanism is executed as follows:
To reduce computational complexity, a Convolution Spatial Reduction (CSR) [wang2021pyramid] is employed, which uses a $k\times k$ depth-wise convolution with a stride of $k$ to downsample the spatial dimensions of $x$ before the linear projection operation. This results in $\widetilde{x}\in{\mathbb{R}^{N_{s}\times C}}$, where $N_{s}=\frac{N}{{{k^{2}}}}$. In our work, we propose a Strip Self-Attention (SSA) module as an innovative token mixing mechanism in the HPB to balance global feature extraction and the computational efficiency of self-attention.
As shown in Fig. 3 (d), the channel dimensions of the query $Q$ and key $K$ are compressed into a single dimension to further minimize computational overhead. Experimental results demonstrate that a strip-shaped configuration for $Q\in{\mathbb{R}^{N\times 1}}$ and $K\in{\mathbb{R}^{N_{s}\times 1}}$ can effectively capture global similarities. The pseudocode for the entire SSA operation is presented in Algorithm 1, followed by a detailed complexity analysis.
Fig. 4 compares the proposed Strip Self-Attention with several notable self-attention variants from prior works. (a) shows the vanilla self-attention used in ViTs [dosovitskiy2020image], which computes global attention via standard dot-product operations. (b) depicts the separable self-attention in MobileViT-v2 [mehta2022separable], which forms a context vector via element-wise interactions between query ($Q$) and key ($K$), then multiplies it with value ($V$) to produce the output. (c) presents the swift self-attention in SwiftFormer [shaker2023swiftformer], an additive variant that relies solely on $Q$ and $K$: the query matrix is reweighted with learnable parameters, pooled into global queries, broadcast, and multiplied element-wise with $K$ to obtain a global context. (d) illustrates the convolutional additive self-attention in CAS-ViT [zhang2024cas], which replaces the core token mixer with a sequence of convolutional spatial attention followed by channel attention, inspired by CBAM [woo2018cbam]. Our Strip Self-Attention differs in two key respects. Firstly, the mechanisms in (b) and (c) operate at the attention-map level using element-wise products rather than full matrix attention, limiting their ability to model true global context. Secondly, CAS-ViT’s [zhang2024cas] purely convolutional path struggles to capture dense token correlations. In contrast, our method jointly reduces redundancy in spatial and channel dimensions, minimizing computational load while preserving global awareness and effectively modeling dense token interactions.
To facilitate clearer understanding, the effective receptive fields (ERFs) of various architectures are distinctly visualized in Fig. 5. Convolution-based methods display highly localized receptive fields, which restrict their ability to capture global context. In contrast, the Vanilla ViT [dosovitskiy2020image] covers all spatial positions, allowing for comprehensive global information extraction, but at the cost of significantly increased computational complexity. Our method achieves a balance between these two extremes. As illustrated, the receptive field in our approach initially focuses on the most critical regions and then progressively expands in a strip-like manner along the four cardinal directions, effectively emulating human visual perception mechanisms.
III-B2 Local Interaction Module (LIM)
where $SE$ denotes the channel attention module [hu2018squeeze].
@@ III-C Complexity Analysis
To demonstrate the efficiency of our proposed SSA, we compared its computational overhead with standard self-attention [dosovitskiy2020image]. Table III presents a detailed computational complexity comparison of the two attention mechanisms.
Specifically, given an input feature $x\in\mathbb{R}^{N\times d}$, the total computational cost of Multi-Head Self-Attention (MHSA), measured in FLOPs, is given by
@@
In the case of our SSA, a spatial reduction is introduced, with $N_{s}=\frac{N}{k^{2}}$, where $k$ represents the kernel size of the spatial reduction convolution.
@@
$\mathcal{O}(\text{SSA})=\frac{1}{k^{2}}Nd^{2}+\frac{h+d}{k^{2}}N^{2}+\left(1+\frac{1}{k^{2}}\right)Ndh.$ (10)
@@
$\mathcal{O}(\text{SSA})\ll\frac{1}{{{k^{2}}}}N{d^{2}}+\frac{{2d}}{{{k^{2}}}}{N^{2}}+\left({1+\frac{1}{{{k^{2}}}}}\right)N{d^{2}}.$ (11)
Let $\mathcal{O}(\text{A})=\frac{1}{{{k^{2}}}}N{d^{2}}+\frac{{2d}}{{{k^{2}}}}{N^{2}}+\left({1+\frac{1}{{{k^{2}}}}}\right)N{d^{2}}$, the complexity can be further simplified as
$\mathcal{O}(\text{A})$ $=\left(1+\frac{2}{k^{2}}\right)Nd^{2}+\frac{2d}{k^{2}}N^{2}$ (12)
$<3Nd^{2}+2dN^{2}.$
@@
TABLE V: Comparison of semantic segmentation performance on the ADE20K dataset [zhou2017scene].
Model Semantic FPN [kirillov2019panoptic]
@@
ResNet-18 [he2016deep] - - 32.9
EMO-1M [zhang2023rethinking] 5 23 34.2
PVT-T [wang2021pyramid] 17 33 35.7
ResNet-50 [he2016deep] 29 46 36.7
S2AFormer-mini (Ours) 6 23 36.7
PoolFormer-S12 [yu2022metaformer] 16 31 37.2
PVTv2-B0 [wang2021pvtv2] 8 25 37.2
EMO-2M [zhang2023rethinking] 6 24 37.3
CASViT-XS [zhang2024cas] 7 24 37.1
S2AFormer-T (Ours) 7 25 38.0
FastViT-SA12 [vasu2023fastvit] 14 29 38.0
ResNet-101 [he2016deep] 48 65 38.8
EfficientFormer-L1 [li2022efficientformer] 16 28 38.9
ResNeXt-101-32x4d [xie2017aggregated] 47 65 39.7
S2AFormer-XS (Ours) 8 26 39.2
LSNet-T [wang2025lsnet] - - 40.1
ResNeXt-101-64x4d [xie2017aggregated] 86 104 40.2
EMO-5M [zhang2023rethinking] 9 26 40.4
PVT-S [wang2021pyramid] 28 45 39.8
RepViT-M1.1 [wang2024repvit] - - 40.6
PoolFormer-S24 [yu2022metaformer] 25 39 40.3
S2AFormer-S (Ours) 12 28 40.8
SwiftFormer-L1 [shaker2023swiftformer] 16 30 41.4
EfficientViM-M4 [lee2025efficientvim] - - 41.3
PoolFormer-S36 [yu2022metaformer] 35 48 42.0
EfficientFormerV2-S2 [li2023rethinking] 16 28 42.4
iFormer-M [zheng2025iformer] - - 42.4
PVTv2-B1 [wang2021pvtv2] 18 34 42.5
MobileMamba-B4 [he2024mobilemamba] 20 - 42.5
PoolFormer-M48 [yu2022metaformer] 77 82 42.7
InceptionNeXt-T [yu2024inceptionnext] 28 44 43.1
FastViT-SA36 [vasu2023fastvit] 34 44 42.9
LSNet-B [wang2025lsnet] - - 43.0
EfficientMod-S [ma2024efficient] 33 - 43.5
S2AFormer-M (Ours) 26 43 43.7
TABLE VI: Comparison of object detection and instance segmentation performance (%) on the COCO val2017 dataset [lin2014microsoft]. FLOPs are tested on images of size $800\times 1280$. Our results, highlighted in bold, demonstrate superior performance with comparative computational overhead.